OpenStudy (anonymous):

What is the simplified form of (15x^7/9y^5) times (45y^2/5x^3)

OpenStudy (anonymous):

\[\frac{15x^7}{9y^5} \times \frac{45y^2}{5x^3}\]

OpenStudy (anonymous):

Do you see any cancellations you can make

OpenStudy (anonymous):

i dont kno about the expoents but u can cancle the 5 out and the 15 would become 3 and u can cancle the nine out and the 45 would be 5

OpenStudy (anonymous):